# The Brownian loop measure on Riemann surfaces and applications to length spectra

### v2

@@ Appendix A Basics on Brownian motion on Riemannian manifolds
We will consider the Brownian motion run at speed $2$, that is $W_{t}=B_{2t}$, and consider $W$ as a random variable in the space of $C^{0}([0,\infty),\mathbb{R}^{n})$ endowed with the topology of uniform convergence on compact subsets of $[0,\infty)$. The advantage of considering the speed $2$ Brownian motion is that its diffusion generator is given by the Laplacian $\Delta=\sum_{k=1}^{n}\partial_{kk}$ (instead of $\Delta/2$). In fact, it is straightforward to check that the distribution of $W_{t}$ starting from $x\in\mathbb{R}^{n}$ is given by
@@
The Brownian motion on a more general complete Riemannian surface $(X,g)$ is the projection of a Brownian motion by a locally isometric universal covering map $\tilde{X}\to X$, where $\tilde{X}=\mathbb{H}$ or $\mathbb{C}=\mathbb{R}^{2}$ or $\mathbb{S}^{2}$, equipped with a conformal metric $g=\mathrm{e}^{2\sigma}(\mathrm{d}x_{1}^{2}+\mathrm{d}x_{2}^{2})$.

### v3

@@ Appendix A Basics on Brownian motion on Riemannian manifolds
We will consider the Brownian motion run at speed $2$, that is $W_{t}=B_{2t}$, and consider $W$ as a random variable in the space of $C^{0}([0,\infty),\mathbb{R}^{n})$ endowed with the topology of uniform convergence on compact subsets of $0,\infty)$. The advantage of considering the speed $2$ Brownian motion is that its diffusion generator is given by the Laplacian $\Delta=\sum_{k=1}^{n}\partial_{kk}$ (instead of $\Delta/2$). In fact, it is straightforward to check that the distribution of $W_{t}$ starting from $x\in\mathbb{R}^{n}$ is given by
@@
The Brownian motion on a more general complete Riemannian surface $(X,g)$ is the projection of a Brownian motion by a locally isometric universal covering map $\tilde{X}\to X$, where $\tilde{X}$$\mathbb{H}$ or $\mathbb{C}=\mathbb{R}^{2}$ or $\mathbb{S}^{2}$, equipped with a conformal metric $g=\mathrm{e}^{2\sigma}(\mathrm{d}x_{1}^{2}+\mathrm{d}x_{2}^{2})$.